OpenStudy (anonymous):

A cellular phone company offers a contract for which the cost c, in dollars, of T minutes of telephoning is given by c=0.25(t-500) +63.95, where it is assumed that t >= 500 minutes. what times will keep costs between $104.20 and $127.20

OpenStudy (anonymous):

down limit is 0.2(t-500)=40.25 at t=701.25 upper limit is 0.2(t-500)=63.25 at t=816.25 sp 701.25<=t<=816.25

OpenStudy (anonymous):

i am not sure i undeerstand

OpenStudy (anonymous):

set C = 104.20 and solve for t: 104.2 = .025(t-500) + 63.95 40.25 = .25(t-500) 161 = t - 500 t = 661 minutes set c = 127.20 127.2 = .25(t-500) + 63.95 63.25 = .25(t-500) 253 = t - 500 t = 753 minutes so 661 minutes =< t =< 753 minutes